OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (anonymous):

y=3x-5

OpenStudy (jedd):

y=3z+d to find d put x=1 , y=-2 -2=3+d d=-5 y=3z-5

OpenStudy (anonymous):

thanks for the explanation!

OpenStudy (jedd):

(: